OpenStudy (anonymous):

Express each repeating decimal as a fraction. 0.6, 0.12, 0.82. Note: there is supposed to be a line over 6, 2, and 82

OpenStudy (anonymous):

- is it 0.6 ????

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

are u supposed to show the working or give answer by shortcut

OpenStudy (anonymous):

I just dont know how to do it. Can you show the work on one and answer the rest so I can check them?

OpenStudy (anonymous):

since there is only one digit repeating here (6) so our aim is to get one 6 before decimal point. - Let p = 0.6 (= 0.6666---- (this goes on endlessly) ---- (1) now to get one 6 before the decimal point multiply both sides by 10 - 10p = 6.6 (= 6.666------ ) -------(2) Subtract (1) from (2) - - 10p - p = 6.6 - 0.6 9p = 6 p = 6/9 p = 2/3 - 2 so 0.6 = ----- 3

OpenStudy (anonymous):

got it?????

OpenStudy (anonymous):

got it. so what is 0.12...2 repeating and 0.82...8and2 repeating

OpenStudy (anonymous):

- now we come to 0.12 (= 0.1222222-------) Here we find that before the repeating digit 2, there is one non-repeating digit i.e.1 so we have to first bring this 1 before the decimal and then 2 once before the decimal point. we proceed as follows: - Let p = 0.12 ------(1) Multiply both sides of (1) by 10 - 10p = 1.2 ------(2) Now multiply both sides of (1) by 100 - 100p = 12.2 ------(3) subtract (2) from (3) - - 100p - 10p = 12.2 - 1.2 90p = 11 p = 11/90 - 11 so 0.12 = ------ 90

OpenStudy (anonymous):

got it????

OpenStudy (anonymous):

got it and what is just the answer for 0.82

OpenStudy (anonymous):

-- 0.82 (= 0.828282.......) Here we have two repeating digits so we have to get one pair of 82 before the decimal point.. -- Let p = 0.82 ------------- (1) multiply both sides of (1) by 100 -- 100p = 82.82 -----------(2) subtracting (1) from (2) -- -- 100p - p = 82 .82 - 0.82 99p = 82 p = 82/99 -- 82 so 0.82 = ------- 99

OpenStudy (anonymous):

is the concept/method clear now???

OpenStudy (anonymous):

yes thank u!

OpenStudy (anonymous):

u r welcome :)) if u wish u can become my fan .....